Anita invested a certain amount of money into a savings account, and invested twice that amount into a mutal fund. If after one year of time the savings account earned 5% interest and the mutual fund earned 8% interest, and total earned from the two investments in that year was $105, how much did Anita initially invest into both?
Answer by checkley77(12569)

.05x+.08*2x=105
.05x+.16x=105
.21x=105
x-105/.21
x=500 was the investment @ 5%
2*500=1,000 is the amount invested @ 8%
proof:
.05*500+.08*1000=105
25+80=105
105=105